# A path-dependent PDE solver based on signature kernels

### v3

@@ 4.2. Consistency
$\sup_{\omega\in\mathcal{X}}\left\lvert\partial u_{m,n}(\omega)\right\rvert\leq C\left\lVert u^{\star}\right\rVert_{\mathcal{H}},\quad\text{for all }\partial\in\mathcal{D}.$
@@
Under condition (i), the same conclusion holds except that $u_{\infty}$ is an element of the closure of $\mathcal{H}$ with respect to the $\mathcal{C}^{\mathcal{N}}$ norm. The universality property of the kernel derived in Lemma 3.17 extends to $\mathcal{C}^{\mathcal{N}}$ by Proposition 4.1 and entails that ¯$\overline{\mathcal{H}}=\mathcal{C}^{\mathcal{N}}(\mathcal{X})$. Indeed, for all $h\in\mathcal{C}^{\mathcal{N}}(\mathcal{X})$ there is a sequence $(h_{k})_{k\in\mathbb{N}}$ in $\mathcal{H}$ that converges to $h$ in $\mathcal{C}_{0}$. Moreover, for all $\partial\in\mathcal{D}$ and $\omega\in\mathcal{X}$, $\partial h_{k}(\omega)=\langle h_{k},\partial\kappa(\omega,\cdot)\rangle_{\mathcal{H}}$ hence $\partial h_{k}$ also converges in $\mathcal{C}_{0}$ towards $\partial h$.
@@ 5. Numerical experiments
In this section, we present numerical experiments benchmarking the signature kernel PPDE solver presented in the previous section against either analytic solutions (if they are available) or classical Monte Carlo solvers. We consider two examples mentioned in subsection 2.3, namely the path-dependent heat equation (2.7)) and the rough Bergomi PPDE (2.9). For both experiments, we consider two types of errors between the predicted prices and true prices, namely the mean squared error (MSE), and the mean absolute error (MAE). The optimal kernel hyperparameters are determined by cross-validation. We will conclude the section with a discussion to compare our kernel approach with recent neural networks techniques for solving PPDEs. Our code is available at https://github.com/crispitagorico/sigppde.
@@ 5.1. Fractional Brownian motion
In this first example, we consider a one dimensional fractional Brownian motion $(\widehat{W}_{t}=\int_{0}^{t}K(t,r)\,\mathrm{d}W_{r})_{t\in\mathbb{T}}$ with Hurst exponent $H\in(0,\frac{1}{2})$. We recall that $\Theta^{t}_{s}=\mathbb{E}[\widehat{W}_{s}|\mathcal{F}_{t}]$ for $s\geq t$. By [54, Theorem 4.1], under appropriate regularity assumptions on the functions $\phi,f$, the conditional expectation
@@
    $f(x)=x,\quad\quad\quad\quad\ \mathbb{E}[f(\widehat{W}_{T})|\mathcal{F}_{t}]=\Theta_{T}^{t}$;
@@
    $f(x)=\mathrm{e}^{\nu x},\quad\quad\quad\ \ \mathbb{E}[f(\widehat{W}_{T})|\mathcal{F}_{t}]=\exp\left(\nu\Theta_{T}^{t}+\frac{\nu^{2}(T-t)^{2H}}{2}\right)$;
@@
    $f(x)=(x-K)_{+},\quad\mathbb{E}[f(\widehat{W}_{T})|\mathcal{F}_{t}]=\frac{(T-t)^{H}}{\sqrt{2\pi}}\exp\left(-\frac{(K-\Theta^{t}_{T})^{2}}{2(T-t)^{2H}}\right)-(K-\Theta^{t}_{T})\vartheta\left(\frac{\Theta^{t}_{T}-K}{(T-t)^{H}}\right)$,
@@
Because analytic prices are available, we limit ourselves to asses the performance of our kernel algorithm to recover these true prices. Collocation points for the kernel method were chosen uniformly on $[0,1]$ for the time variable and sampled from the process $\Theta^{t}$ for the path variable.
@@ 5.2. Rough Bergomi
for independent Brownian motions $(W^{1},W^{2})$. The hybrid scheme of [8] is used for efficient, $\mathcal{O}(\ell\log\ell)$, simulation of the Volterra process $\widehat{W}$ where $\ell$ is the length of the simulated paths. For the experiments we take the same parameters as in the repository [^2] https://github.com/ryanmccrickerd/rough_bergomi/tree/master., namely $T=1$, $\xi=0.055$, $\eta=1.9$, $\rho=-0.9$. We set the ground truth to be the Monte Carlo prices obtained with $100000$ sample paths. We recall that the value function $u$ defined in subsection 2.3.3 (and dubbed $u_{2}$ in the introduction) solves the PPDE
@@
if $s>t$ and 00 otherwise, the value function $u$ has the following probabilistic representation
@@
since $\Theta^{t}$ is an $\mathcal{F}_{t}$ measurable process and $I^{t}$ is independent from $\mathcal{F}_{t}$. Collocation points for the kernel method were sampled uniformly at random on $[0,1]$ for the time variable, uniformly at random on $[x_{min},x_{max}]$ for the price variables, where $x_{min},x_{max}$ were chosen using ground truth prices, and sampled from the process $\Theta^{t}$ for the path variable.
$H=0.1$ – Strike $=0.1$
@@
We consider two values of $H=0.1$ and $H=0.3$ and two strike values $K=0.1$ and $K=1.0$. We take the same number of sample paths and collocation points for Monte Carlo and the signature kernel solver respectively, and consider three different increasing values, namely $20$, $100$ and $500$. As it can be observed from the tables, the numerical performance in terms of MSE and MAE of the two algorithms are comparable across the different settings.
@@
Nevertheless, the PDE approach provides by design advantages over Monte Carlo methods. Indeedn, finite-dimension greeks are prone to instability, let alone pathwise ones, while all derivatives of the PPDE solution appear from (the linear combination of) standard PDEs.
@@ 5.3. Comparison with neural PDE solvers
As antiticipated, we will conclude this section with some remarks aimed at comparing our kernel approach to the numerous neural network techniques to solve PPDE which have been proposed in the literature in recent years. Neural PDE solvers such as DGM [51] and PINNs [43] essentially parameterise the solution of the PDE as a neural network which is then trained using the PDE and its boundary conditions as loss function. One drawback of these approaches in the path-dependent setting [45, 50, 31] is that they require a time-discretisation of the solution. The time grids over which solutions and derivatives are evaluated during training and testing need to agree, usually followed by an arbitrary interpolation. On the contrary, our kernel approach operates directly on continuous paths, and it is therefore mesh-free. Furthermore, derivatives can be accessed directly by differentiating kernels, with no need to prematurely discretise in time. We note that concomitant to this paper is [23], which uses a neural rough differential equation (Neural RDE) model [40, 48] to parameterise the solution of a PPDE and is therefore also mesh-free. Another drawback of neural PDE solvers is that their theoretical analysis is often limited to density or universal approximation results; showing the existence of a network of a requisite size achieving a certain error rate, without guarantees whether this network is computable in practice. In fact the optimisation, done with variants of stochastic gradient descent, is never convex so there is no guarantee to attain a global minimum. Our kernel approach boils down to a convex finite dimensional optimisation, guaranteed to attain the unique global minimum, and that is consistent with the original problem when the number of collocation points is sent to infinity.
@@ A.1. Differentiability and continuity estimates for the signature
$\quad+\int_{t}^{t^{\prime}}\sum_{i,j=1}^{e}\left\lVert\partial^{2}_{x^{ij}}g(\gamma_{u},\cdot)\right\rVert_{\mathcal{H}_{g}}\Big{(}\left\lvert\eta_{u}\right\rvert\left\lvert\,\mathrm{d}{\bar{\eta}}_{u}\right\rvert+\left\lvert{\bar{\eta}}_{u}\right\rvert\left\lvert\,\mathrm{d}\eta_{u}\right\rvert\Big{)}$

### v4

@@ 4.2. Consistency
$\sup_{\omega\in\mathcal{X}}\left\lvert\partial u_{m,n}(\omega)\right\rvert\leq C\left\lVert u^{\star}\right\rVert_{\mathcal{H}},\quad\text{for all }\partial\in\mathscr{D}.$
@@
Under condition (i), the same conclusion holds except that $u_{\infty}$ is an element of the closure of $\mathcal{H}$ with respect to the $\mathcal{C}^{\mathcal{N}}$ norm. The universality property of the kernel derived in Lemma 3.17 extends to $\mathcal{C}^{\mathcal{N}}$ by Proposition 4.1 and entails that $\overline{\mathcal{H}}=\mathcal{C}^{\mathcal{N}}(\mathcal{X})$. Indeed, for all $h\in\mathcal{C}^{\mathcal{N}}(\mathcal{X})$ there is a sequence $(h_{k})_{k\in\mathbb{N}}$ in $\mathcal{H}$ that converges to $h$ in supremum norm. Moreover, for all $\partial\in\mathscr{D}$ and $\omega\in\mathcal{X}$, $\partial h_{k}(\omega)=\langle h_{k},\partial\kappa(\omega,\cdot)\rangle_{\mathcal{H}}$ hence $\partial h_{k}$ also converges in supremum norm towards $\partial h$.
@@ 5. Numerical experiments
In this section, we present numerical experiments benchmarking our signature kernel PPDE solver formulated as the optimal recovery problem (2.17), whose exact solution is given by (2.18) with coefficients obtained by solving the linear systems (2.19)–(2.20) against either analytic solutions (when available) or classical Monte Carlo solvers. We consider two examples mentioned in subsection 2.3, namely the path-dependent heat equation (2.10)) and the rough Bergomi PPDE (2.12). For both experiments, we consider two types of errors between the predicted prices and true prices, namely the mean squared error (MSE), and the mean absolute error (MAE). The optimal kernel hyperparameters are determined by cross-validation. We will conclude the section with a discussion to compare our kernel approach with recent neural networks techniques for solving PPDEs. Our code is available at https://github.com/crispitagorico/sigppde.
@@ 5.1. Fractional Brownian motion
In this first example presented in Section 2.3.1, we consider a one dimensional fractional Brownian motion $(\widehat{W}_{t}=\int_{0}^{t}K(t,r)\,\mathrm{d}W_{r})_{t\in\mathbb{T}}$ with Hurst exponent $H\in(0,\frac{1}{2})$. We recall that $\Theta^{t}_{s}=\mathbb{E}[\widehat{W}_{s}|\mathcal{F}_{t}]$ for $s\geq t$. By [56, Theorem 4.1], under appropriate regularity assumptions on the functions $\phi,f$, the conditional expectation
@@
    $\phi(x)=x,\quad\quad\quad\quad\ \mathbb{E}[\phi(\widehat{W}_{T})|\mathcal{F}_{t}]=\Theta_{T}^{t}$;
@@
    $\phi(x)=\mathrm{e}^{\nu x},\quad\quad\quad\ \ \mathbb{E}[\phi(\widehat{W}_{T})|\mathcal{F}_{t}]=\exp\left(\nu\Theta_{T}^{t}+\frac{\nu^{2}(T-t)^{2H}}{2}\right)$;
@@
    $\phi(x)=(x-K)_{+},\quad\mathbb{E}[\phi(\widehat{W}_{T})|\mathcal{F}_{t}]=\frac{(T-t)^{H}}{\sqrt{2\pi}}\exp\left(-\frac{(K-\Theta^{t}_{T})^{2}}{2(T-t)^{2H}}\right)-(K-\Theta^{t}_{T})\vartheta\left(\frac{\Theta^{t}_{T}-K}{(T-t)^{H}}\right)$,
@@
Because analytic prices are available, we limit ourselves to assess the performance of our kernel algorithm to recover these true prices. Collocation points for the kernel method were chosen uniformly on $[0,1]$ for the time variable and sampled from the process $\Theta^{t}$ for the path variable.
@@ 5.2. Rough Bergomi
for independent Brownian motions $(W^{1},W^{2})$. The hybrid scheme of [7] is used for efficient, $\mathcal{O}(\ell\log\ell)$, simulation of the Volterra process $\widehat{W}$ where $\ell$ is the length of the simulated paths. For the experiments we take the same parameters as in the repository [^2] https://github.com/ryanmccrickerd/rough_bergomi/tree/master., namely $T=1$, $\xi=0.055$, $\eta=1.9$, $\rho=-0.9$. We set the ground truth to be the Monte Carlo prices obtained with $100000$ sample paths. We recall that the value function $u$ defined in subsection 2.3.3 solves the PPDE
@@
if $s>t$ and 0 otherwise, the value function $u$ has the following probabilistic representation
@@
since $\Theta^{t}$ is an $\mathcal{F}_{t}$ measurable process and $I^{t}$ is independent from $\mathcal{F}_{t}$. Collocation points for the kernel method were sampled uniformly at random on $[0,1]$ for the time variable, uniformly at random on $x_{min},x_{max}]$ for the price variables, where $x_{min},x_{max}$ were chosen using ground truth prices, and sampled from the process $\Theta^{t}$ for the path variable.
$H=0.1$ – Strike =0.1=0.1
@@
We consider two values of $H=0.1$ and $H=0.3$ and two strike values $K=0.1$ and $K=1.0$. We take the same number of sample paths and collocation points for Monte Carlo and the signature kernel solver respectively, and consider three different increasing values, namely $20$, $100$ and $500$. As it can be observed from the tables, the numerical performance in terms of MSE and MAE of the two algorithms are comparable across the different settings. As shown in the tables above, the relative performance depends on the roughness of the input paths (governed by the Hurst parameter $H$) and on the strike of the option. Hence, the comparison is not one-sided.
@@
Nevertheless, the PDE approach provides by design advantages over Monte Carlo methods. Indeed, finite-dimension greeks are prone to instability, let alone pathwise ones, while all derivatives of the PPDE solution appear from (the linear combination of) standard PDEs.
@@ 5.3. Comparison with neural PDE solvers
As anticipated, we will conclude this section with some remarks aimed at comparing our kernel approach to the numerous neural network techniques to solve PPDE which have been proposed in the literature in recent years. Neural PDE solvers such as DGM [53] and PINNs [43] essentially parameterise the solution of the PDE as a neural network which is then trained using the PDE and its boundary conditions as loss function. One drawback of these approaches in the path-dependent setting [47, 52, 31] is that they require a time-discretisation of the solution. The time grids over which solutions and derivatives are evaluated during training and testing need to agree, usually followed by an arbitrary interpolation. On the contrary, our kernel approach operates directly on continuous paths, and it is therefore mesh-free. Furthermore, derivatives can be accessed directly by differentiating kernels, with no need to prematurely discretise in time. We note that concomitant to this paper is [23], which uses a neural rough differential equation (Neural RDE) model [40, 50] to parameterise the solution of a PPDE and is therefore also mesh-free. Another drawback of neural PDE solvers is that their theoretical analysis is often limited to density or universal approximation results; showing the existence of a network of a requisite size achieving a certain error rate, without guarantees whether this network is computable in practice. In fact the optimisation, done with variants of stochastic gradient descent, is never convex so there is no guarantee to attain a global minimum. Our kernel approach boils down to a convex finite-dimensional optimisation, guaranteed to attain the unique global minimum, and that is consistent with the original problem when the number of collocation points is sent to infinity.
@@ A.1. Differentiability and continuity estimates for the signature
[…]